OpenStudy (anonymous):

Please Help this is killing me right now. I know how to solve this word problem once it is set up, but I just don't know how to set it up. The radiator of a car needs to contain 40 liters of 40% anti freeze. The radiator now contains 40 liters of 20% antifreeze. How many liters of this solution must be drained and replaced with 100% antifreeze to get the desired strength?

OpenStudy (amistre64):

amount(%) is the key to this stuff: x+y = 40; y = 40-x ................................... replace the y with its 'value' from above and solve for x x(.2) + y(1) = 40(.4)

OpenStudy (amistre64):

makes sense?

OpenStudy (anonymous):

so 30 liters? 30 liters must stay so 10 liters need to be drained?

OpenStudy (amistre64):

dunno; let me do the math :)

OpenStudy (anonymous):

okay haha

OpenStudy (anonymous):

This problem is haunting my life right now. It's the devil...

OpenStudy (amistre64):

.2x + 40 - x = 16 .2x - x = 16-40 -.8x = -24 x = 24/.8 = 240/8 = 30 there should be 30 liters of 20% so we need to drain 10 liters and top it with 100%

OpenStudy (amistre64):

lets check that :)

OpenStudy (amistre64):

30(.2) + 10(1) = 40(.4) 6 + 10 = 16 16 = 16....we are goood

OpenStudy (anonymous):

OMG! i did that, but for some reason i thought it was wrong.... lame thanks dude!

OpenStudy (amistre64):

:) youre welcome

OpenStudy (preetha):

good work Amistre64. Nice set up of the equation.